# Benchmarking News Recommendation in the Era of Green AI

### 07 Mar 2024 v1

@@ Abstract.
Over recent years, news recommender systems have gained significant attention in both academia and industry, emphasizing the need for a standardized benchmark to evaluate and compare the performance of these systems. Concurrently, Green AI advocates for reducing the energy consumption and environmental impact of machine learning. To address these concerns, we introduce the first Green AI benchmarking framework for news recommendation, known as GreenRec [^1] GreenRec is published under the Legommenders benchmark, a modular framework for recommender systems. We have released the source code and data at https://github.com/Jyonn/Legommenders for reproducibility., and propose a metric for assessing the tradeoff between recommendation accuracy and efficiency. Our benchmark encompasses 30 base models and their variants, covering traditional end-to-end training paradigms as well as our proposed efficient only-encode-once (OLEO) paradigm. Through experiments consuming 2000 GPU hours, we observe that the OLEO paradigm achieves competitive accuracy compared to state-of-the-art end-to-end paradigms and delivers up to a 2992% improvement in sustainability metrics.
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3651472isbn: 979-8-4007-0172-6/24/05
@@ 1. Introduction
News recommender systems have become a crucial tool in the digital age, helping users discover relevant news articles and stay informed amidst the ever-increasing number of sources available. To integrate news content knowledge into recommender systems, content encoders are employed to capture semantic news representations. Various approaches have been explored in these efforts, including simple content encoders represented by CNN (Kim, 2014) and Attention-based (Vaswani et al., 2017) models (An et al., 2019; Wu et al., 2019a, c), as well as emerging content encoders based on pretrained language models (Zhang et al., 2021; Liu et al., 2022).
@@ 2.1. Basics of News Recommender Systems
News recommender systems are devised to predict the click probability of a user over a piece of given news article. Typically, a general news recommender system consists of three main components: content encoder, history encoder, and interaction module. As depicted in Figure 1 (a), the recommender accepts a news-user pair as input and yields the click probability. The content encoder is responsible for capturing the semantic meaning and contextual information of the news articles, while the history encoder processes the news history vectors, capturing the preferences, interests, and browsing patterns of users. Finally, the interaction module (e.g., a dot product operation) facilitates the interaction between the candidate news vector and user vector, ultimately calculating the click probability or relevance score for the given news-user pair.
@@ 2.2. Training Paradigms and Complexity
Traditional models (Wu et al., 2019b, a, c, 2021) typically follow such end-to-end training paradigm, as displayed in Figure 1 (a), where the three components are jointly trained under the supervision of the click label with the recommendation task.
While the recommendation accuracy improves through the adoption of these more sophisticated component designs, such conventional paradigm has gradually revealed their inherent inefficiency. The primary reason behind lies in the redundant processing within the content encoder: in the training process, the same news articles are excessively encoded due to the frequent appearance of users. With $N$ news articles, $M$ users, $K$ user-news training pairs, and an average user historical sequence length of $L$, a single end-to-end training process involves processing $M\times L$ news articles by the content encoder. Given time complexities $t_{c}$, $t_{u}$, and $t_{i}$ for the content encoder, user encoder, and interaction module, respectively, the end-to-end training time complexity is approximately $O(M\times L\times t_{c}+K\times(t_{u}+t_{i}))$, where $M\times L>>N,orK$. The introduction of a large-scale content encoder (Zhang et al., 2021; Wu et al., 2021), with $t_{c}>>t_{u}+t_{i}$, simplifies the time complexity to $O(M\times L\times t_{c})$. A more specific example is within the MIND dataset, where, on average, each news article will be encoded approximately 1818 (i.e., $M\times L/N$) times during a single training epoch.
Inspired by the work of Liu et al. (2022), we introduce an only-encode-once (OLEO) training paradigm, as demonstrated in Figure 1 (b), where the content encoder undergoes an initial pretraining phase using news content in a self-supervised manner. Once the pretraining is complete, the content encoder proceeds to extract and cache the content vectors (i.e., news representations). Subsequently, the other two modules, namely the history encoder and interaction module, are trained for the recommendation task. All the news representations can be easily and quickly accessed through a vector lookup table. In this paradigm, the time complexity for content encoder pretraining is $O(N\times t_{c})$, which can be treated as a common overhead used across different base models, while the time complexity for the recommender is $O(K\times(t_{u}+t_{i}))$, which is mush smaller than that of the end-to-end paradigm.
@@ 2.4. Proposed GreenRec Benchmark
Usage TorchRec DeepCTR DeepRec RecBole FuxiCTR BARS Ours
Data Preprocessing
Content-aware CTR $\times$ $\times$ $\times$
NewsRec Methods $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Decoupled Training $\times$ $\times$ $\times$
End-to-end Training $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Green AI Evaluation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@ 3.1. Comparison with Previous Benchmarks
In recent years, the issues of non-reproducibility and unfair comparisons have garnered increasing attention within the recommendation community. A series of studies (Ivchenko et al., 2022; Shen, 2017; Zhang et al., 2019; Zhao et al., 2022; Zhu et al., 2021, 2022) have conducted comprehensive experiments and analyses, advocating for a unified benchmark construction. As shown in Table 1, we compare our news recommendation benchmark GreenRec with these existing efforts. It is evident that these studies primarily target in the general recommendation domain, particularly emphasizing aspects such as CTR prediction and collaborative filtering. Nevertheless, a standardized benchmarking pipeline tailored to the news recommendation domain and the environmental evaluation remains conspicuously absent. Hence, our GreenRec is the first attempt on benchmarking the news recommenders, encompassing two training paradigms and five variants for each base model.
@@ 3.2. Dataset and Data Splitting
MIND-small MIND-large
##\## News 65,238 104,151
##\## Users 94,057 750,434
##\## Interactions 347,727 3,958,501
##\## Samples 8,381,093 95,447,571
Density 0.1366% 0.1221%
@@
NewsRec CTR NewsRec CTR
Method NAML LSTUR NRMS BST DCN DIN NAML LSTUR NRMS BST DCN DIN
@@
Due to the fact that the MIND-small dataset only provides training and validation sets, many studies have failed to specify their exact partitioning during training, and in some cases, they have directly reported results on the validation set, which leads to an unfair comparison among different models. Furthermore, the MIND-large dataset only offers online testing, which is not conducive to offline environment assessment. To address these issues, we have restructured both datasets by splitting the validation sets of both datasets in a 1:1 ratio to create new validation and testing sets.
@@ Acknowledgement
Qijiong Liu would like to thank Dr. Guanliang Chen and Mr. Haoxiang Shi for their brainstorming on our paper. Additionally, we would like to thank anonymous reviewers for their feedback.

### 15 Mar 2024 v2

@@ Abstract.
Over recent years, news recommender systems have gained significant attention in both academia and industry, emphasizing the need for a standardized benchmark to evaluate and compare the performance of these systems. Concurrently, Green AI advocates for reducing the energy consumption and environmental impact of machine learning. To address these concerns, we introduce the first Green AI benchmarking framework for news recommendation, known as GreenRec [^1] GreenRec has been published as part of the Legommenders benchmark, which is a modular framework for recommender systems. The source code and data for GreenRec can be found at https://github.com/Jyonn/Legommenders., and propose a metric for assessing the tradeoff between recommendation accuracy and efficiency. Our benchmark encompasses 30 base models and their variants, covering traditional end-to-end training paradigms as well as our proposed efficient only-encode-once (OLEO) paradigm. Through experiments consuming 2000 GPU hours, we observe that the OLEO paradigm achieves competitive accuracy compared to state-of-the-art end-to-end paradigms and delivers up to a 2992% improvement in sustainability metrics.
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3651472isbn: 979-8-4007-0172-6/24/05ccs: Information systems Recommender systemsccs: Information systems Data mining
@@ 1. Introduction
News recommender systems have become a crucial tool in the digital age, helping users discover relevant news articles and stay informed amidst the ever-increasing number of sources available. To integrate news content knowledge into recommender systems, content encoders are employed to capture semantic news representations. Various approaches have been explored in these efforts, including simple content encoders represented by CNN (Kim, 2014) and Attention-based (Vaswani et al., 2017) models (An et al., 2019; Wu et al., 2019a, c; Li et al., 2022), as well as emerging content encoders based on pretrained language models (Zhang et al., 2021; Liu et al., 2022).
@@ 2.1. Basics of News Recommender Systems
News recommender systems are devised to predict the click probability of a user over a piece of given news article. Typically, a general news recommender system consists of three main components: content encoder, history encoder, and interaction module. As depicted in Figure 1 (a), the recommender accepts a news-user pair as input and yields the click probability. The content encoder is responsible for capturing the semantic meaning and contextual information of the news articles, while the history encoder processes the news history vectors, capturing the preferences, interests, and browsing patterns of users. Finally, the interaction module facilitates the interaction between the candidate news vector and user vector to calculate the click probability for the given news-user pair.
@@ 2.2. Training Paradigms and Complexity
Traditional models (Wu et al., 2019b, a, c, 2021) typically follow such end-to-end training paradigm, as displayed in Figure 1 (a), where the three components are are trained together, guided by click-through data.
However, this traditional approach has shown limitations, particularly in terms of efficiency. A significant issue is the repetitive work done by the content encoder due to the repeated encoding of the same news articles, a consequence of their frequent interaction with various users. For instance, if there are $N$ news articles, $M$ users, and $K$ instances of user-news interactions, with an average historical sequence of $L$ interactions per user, the content encoder ends up processing $M\times L$ articles. The overall time required for training can be roughly calculated as $O(M\times L\times t_{c}+K\times(t_{u}+t_{i}))$, highlighting a particularly heavy load when $M\times L$ significantly exceeds either $N$ or $K$, especially with advanced content encoders (Zhang et al., 2021; Wu et al., 2021) that have a much higher time complexity compared to other modules. For example, in the MIND dataset scenario, a single news article is encoded on average about 1818 times in one epoch of training, indicating a substantial amount of redundancy.
To address these inefficiencies, we draw inspiration from a new training approach known as the only-encode-once paradigm, as depicted in Figure 1 (b). This approach involves initially pretraining the content encoder using the news content in a self-supervised manner. After this phase, the content vectors are extracted and stored, allowing the history encoder and interaction module to be trained separately with much quicker access to the news representations via a lookup table. The pretraining phase for the content encoder has a time complexity of $O(N\times t_{c})$, considered a fixed cost across various models. Meanwhile, the training of the recommendation system itself becomes significantly more efficient, with a time complexity of $O(K\times(t_{u}+t_{i}))$, marking a drastic reduction from the traditional method.
Usage TorchRec DeepCTR DeepRec RecBole FuxiCTR BARS Ours
Data Preprocessing
Content-aware CTR $\times$ $\times$ $\times$
NewsRec Methods $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Decoupled Training $\times$ $\times$ $\times$
End-to-end Training $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Green AI Evaluation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@ 3.1. Comparison with Previous Benchmarks
In recent years, the issues of non-reproducibility and unfair comparisons have garnered increasing attention within the recommendation community. A series of studies (Ivchenko et al., 2022; Shen, 2017; Zhang et al., 2019; Zhao et al., 2022; Zhu et al., 2021, 2022) have conducted comprehensive experiments and analyses, advocating for a unified benchmark construction. As shown in Table 1, we compare our news recommendation benchmark GreenRec with these existing efforts. It is evident that these studies primarily target in the general recommendation domain, particularly emphasizing aspects such as CTR prediction and collaborative filtering. Nevertheless, a standardized benchmarking pipeline tailored to the news recommendation domain and the environmental evaluation remains conspicuously absent. Hence, our GreenRec is the first attempt on benchmarking the news recommenders in the era of Green AI.
@@ 3.2. Dataset and Data Splitting
##\## News ##\## Users ##\## Interactions ##\## Samples Density
MIND-small 65,238 94,057 347,727 8,381,093 0.1366%
MIND-large 104,151 750,434 3,958,501 95,447,571 0.1221%
@@
Method Matching CTR Matching CTR
NAML LSTUR NRMS BST DCN DIN NAML LSTUR NRMS BST DCN DIN
@@
Due to the fact that the MIND-small dataset only provides training and validation sets, many studies have failed to specify their exact partitioning during training, and in some cases, they have directly reported results on the validation set, which leads to an unfair comparison among different models. Furthermore, the MIND-large dataset only offers online testing, which is not conducive to offline environment assessment. To address these issues, we have restructured both datasets by splitting the validation sets of both datasets in a 1:1 ratio to create new validation and testing sets. The scripts can be found in our released repository.
@@ Acknowledgement
Qijiong Liu would like to thank Dr. Guanliang Chen and Mr. Haoxiang Shi for their insightful input during the brainstorming process of our paper. Additionally, the authors would like to thank the anonymous reviewers for their valuable feedback.